OpenStudy (anonymous):

does any one know the answers to Unit 2 Lesson 4 Algebra 1 A part 2 for any one who is in connections Academy if u know the answers please message me.

OpenStudy (anonymous):

i really need these answers to bring my grade up if u know them please message them to me

OpenStudy (whpalmer4):

You'll get banned from OpenStudy for trading answers like that. Also, you won't be able to do the problems by yourself in the future, and what you are doing in Algebra 1 is the foundation to nearly everything else. Better that you learn how to do the problems! Put one up, and I'll show you how.

OpenStudy (anonymous):

ok thanks

OpenStudy (whpalmer4):

So, what's the first problem?

OpenStudy (anonymous):

1.Write an equation in point-slope form for the line through the given point with the given slope (8,3);m=6 a.y+3=6(x-8) b.y-3=6(x-8) c.y-3=6(x+8) d.y+3=6x+8

OpenStudy (whpalmer4):

do you know what point-slope form is?

OpenStudy (anonymous):

a little im not that good in algebra

OpenStudy (whpalmer4):

it lets us write the equation for a line if we know the slope \(m\) and any point on the line \((x_0,y_0)\): \[y-y_0 = m(x-x_0)\] so we have \(m=6\) and our point is \((x_0,y_0) = (8,3)\) Good so far?

OpenStudy (anonymous):

i think so

OpenStudy (mathmale):

It's extremely important that you learn to recognize and to use these formulas. I strongly suggest you make up a list of the formulas you've used in solving problems and refer to it often for review and practice. Please, let's ditch the "I'm not that good at algebra" comment; it doesn't help you at all. Good luck.

OpenStudy (mathmale):

All the data you need are already given: the given point with the given slope (8,3);m=6 and whpalmer has given you the appropriate equation. What next?

OpenStudy (anonymous):

i know you have to plug in the numbers to get the equation but i get confused when doing that

OpenStudy (whpalmer4):

Okay, so let's plug in those numbers! first, plug in \(m=6\): \[y - y_0 = 6(x-x_0)\]next, plug in \(x_0 = 8\) \[y-y_0 = 6(x-8)\]finally, plug in \(y_0 = 3\) \[y-3=6(x-8)\] Does that makes sense?

OpenStudy (anonymous):

a little

OpenStudy (mathmale):

whpalmer has taken all of the data and substituted it into the appropriate equation for the straight line. What do you need to do next?

OpenStudy (whpalmer4):

before we do that, is there anything that isn't clear about what I did? do you see why each number went where it did?

OpenStudy (whpalmer4):

you do have to remember that coordinates are listed in the order (x,y) (alphabetical order, as it turns out)

OpenStudy (anonymous):

i remember that but oh i see i think i get it now

OpenStudy (mathmale):

This discussion will prove to be much more valuable to you in the long run if you'd please take notes on it and refer to those notes often later on when you need to apply this procedure.

OpenStudy (whpalmer4):

So, we have \[y-3=6(x-8)\]Does that match any of our answer choices? \[a.y+3=6(x-8)\] \[b.y-3=6(x-8)\] \[c.y-3=6(x+8)\] \[d.y+3=6x+8\]

OpenStudy (whpalmer4):

(I'm having some connectivity issues with OpenStudy, sorry about the slow response!)